Suppose 32% of college students take PE, 52% of college students are female, and 25% of college PE students are female.

(a) What percentage of college students are female and don't take PE?

(b) What percentage of male college students don't take PE?
Answer #1

let probability of female and taken PE are P(F) and P(PE)

therefore P(PE)=32% ; P(F)=52% ; P(F|PE)=25%

a) percentage of college students are female and don't take PE =P(F)-P(F n E)= =P(F)-P(PE)*P(F|PE)=52%-32%*25%=44 %

b)

percentage of male college students don't take PE =P(Fc)-P(PE)*P(Fc|PE)=(100-52)-32*(100-25)% =24%